Find the laplace of cosat where a is a constant


1
Expert's answer
2021-07-16T08:18:40-0400

By definition of a Laplace transform we have

"\mathcal{L}(\cos at)(s)=f(s)=\int_0^{+\infty} e^{-st}\cos(at)\: dt"

We will apply the integration by parts :

"f(s)= \frac{1}{a}\left(\left[e^{-st}\sin(at)\right]^{+\infty}_0+\int_0^{+\infty}se^{-st}\sin(at)\: dt \right)"

The first bracket is 0, as "e^{-st}\to0 \text{ when }t\to +\infty" and "\sin(at)_{t=0}=0". We then apply the integration by parts to the integral :

"f(s)=\left[-\frac{s}{a^2}e^{-st} \cos(at)\right]^{+\infty}_0-\frac{s^2}{a^2} \int_0^{+\infty} e^{-st} \cos(at) \: dt"

The first bracket gives "\frac{s}{a^2}" and the integral is just "f(s)", so we have :

"f(s)=\frac{s}{a^2}-\frac{s^2}{a^2}f(s)"

"\frac{a^2+s^2}{a^2}f(s)=\frac{s}{a^2}"

And therefore, the final answer is "f(s)=\frac{s}{s^2+a^2}" (defined for "s>0").